OpenStudy (anonymous):

5 1/5 divided by 1 3/10? can someone show me how to solve this ?

OpenStudy (amistre64):

id say divide ...

OpenStudy (amistre64):

if you turn them into improper fractions; itll be easier

OpenStudy (anonymous):

i know that , but once they are improper what then

OpenStudy (amistre64):

well the memory schtic is to flip the bottom over the top and multiply across

OpenStudy (amistre64):

but what you really do is just multiply top and bottom by the reciprocal of the bottom

OpenStudy (amistre64):

\[\frac{\cfrac{a}{b}}{\cfrac{c}{d}}*\frac{\cfrac{d}{c}}{\cfrac{d}{c}}=\frac{\cfrac{ad}{bc}}{\cfrac{cd}{cd}}\implies\frac{ad}{bc}\]

OpenStudy (amistre64):

as you can see, the bottom goes to one and becomes rather trivial

OpenStudy (anonymous):

yeah i see it now , just wanted to make i was showing my son the right way to do it